OpenStudy (linn99123):

NEED HELP FAST WILL GIVE MEDAL AND FAN Jethro is planting flowers in his garden. He wants the ratio of tulips to roses to be 7 to 2. Jethro wants to plant a total of 90 flowers. How many tulips should he plant?

OpenStudy (anonymous):

total number of flowers is A + B A/B = 6/2 = 3 --> A = 3B So the total number of flowers = A + B = 3B + B = 4B = 96 --> B = 24 --> A = 3 * 24 = 72

OpenStudy (anonymous):

Did you just copy and paste that?

OpenStudy (anonymous):

*ehm ehm* http://answers.yahoo.com/question/index?qid=20111223105346AAC68FY

OpenStudy (anonymous):

Lol...

OpenStudy (linn99123):

okat then -.-

OpenStudy (anonymous):

next time reference it :P

OpenStudy (linn99123):

dont need peoples smart comments on here so buzz off cause i aint dealing with it

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

LALALALAALA

OpenStudy (linn99123):

not u hun ur my friend lol i meant the one with the werid name typing xD

OpenStudy (anonymous):

oh lol xD sorry

OpenStudy (linn99123):

nah its cool :]

OpenStudy (anonymous):

O.o weird? You're weird you little freak. =.=

OpenStudy (linn99123):

i is a freak actually XD

OpenStudy (anonymous):

>.<

OpenStudy (linn99123):

lol ttyl annas dis chat doe be gettn spicy XD

OpenStudy (mrnood):

if the number of tulips is 72, then the number of roses is 18 the ration is 4:1 - so the above answer is wrong Linn - try working it out - here's a tip 7/9 and 2/9 have a ratio of 7:2

OpenStudy (anonymous):

burnnnnn